y+3=1/4(x+4) thank you
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

I take it you have to solve for x?:
y+3 = 1/4(x+4) Lets rewrite this:
y+3 = (x+4)/4 Since I want to solve for x, I'll move the x to the left side:
(x+4)/4 = y+3 Now multiply both sides by 4:
x+4 = 4y+12 Finally, subtract 4 from both sides:
x = 4y+8

Answer by MathTherapy(10552)


y+3=1/4(x+4) thank you
What is it you wish to do?